OpenStudy (anonymous):

What is an estimate of the solution of the equation 6n + 3 = 2 explain,please.

OpenStudy (anonymous):

\[6n + 3 = 2\] Subtract 3 from both sides \[6n + 3 - 3= 2 - 3\] You get \[6n = -1\] Divide both sides by 6 \[n = -1\6] And that's your answer :)